OpenStudy (anonymous):

Which of the following points lies on the graph of the function y = 2(3)^x a. 2, 0) b. (1, 3) c. (1, 4) d. (0, 2)

OpenStudy (anonymous):

D. (0, 2) Try this graphing calculator: https://www.desmos.com/calculator

OpenStudy (anonymous):

It helps a lot when you want to check your answer

OpenStudy (anonymous):

oops \[2\times 3^0=2\times 1=2\]

OpenStudy (anonymous):

i never learned how to do this though :/ . can you teach me the step by step?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\((a,b)\) is a point on the graph of \(y=f(x)\) means \(f(a)=b\)

OpenStudy (anonymous):

take the \(x\) in the expression \(2\times 3^x\) and replace it by the first coordinate in the ordered pair

OpenStudy (anonymous):

for example in the pair \((2, 0)\) if you replace \(x\) by \(2\) you get \(y=2\times 3^2=2\times 9=18\) and therefore \((2,0)\) is not on the graph, on the other hand \((2,18)\) is on the graph

OpenStudy (anonymous):

yes. is that all of the steps?